OpenStudy (anthonyym):

A person weighing 900 N is standing on snowshoes. Each snowshoe has area 2500 cm^2. What is the pressure exerted by each individual snow shoe on the snow?

OpenStudy (anthonyym):

A) 0.18 N/m^2 B) 0.36 N/m^2 C) 1800 N/m^2 D) 3600 N/m^2 Is it 3600 N/m^2?

OpenStudy (harman.singh):

I have not studied this topic yet but after searching your question on google, 3600 seems to be the correct answer. F = 900N A = 2500cm^2 = 0.25m^3 P=F/A P=900/0.25 P=3600 Source: https://www.physicsforums.com/threads/pressure-force-area-cannot-arrive-at-correct-answer.503230/

OpenStudy (anthonyym):

Yes that's what I think too but the answer key says it is C) 1800 N/m^2, just like many answer keys that I googled online.

OpenStudy (sooobored):

a person has 2 legs, hence the weight of 900N should be divided among the 2 legs evenly so 450N per leg or per snowshoe you know the area of a single snow shoe is 2500cm^2 so you can find pressure \[P=F/A\]

OpenStudy (harman.singh):

Ah okay. This one is better: https://answers.yahoo.com/question/index?qid=20111027101212AArVEyq There are two snowshoes. So each snowshoe apply a pressure of 3600/2

OpenStudy (sooobored):

pressure exerted by a single snowshoe